Find the deBroglie wave length of (a)an electron whose speed is1.0×108m/s,and (b)an electron whose speed is2.0×108m/s.


1
Expert's answer
2020-05-11T20:06:20-0400

By definition, the deBroglie wavelength is given by:

"\lambda = \dfrac{h}{p}" , where "h = 6.6\cdot 10^{-34} J\cdot s" is the Planck constant and "p" is the momentum of the electron. In the relativistic case:

"p = \dfrac{mv}{\sqrt{1-\frac{v^2}{c^2}}}" , where "m = 9.1\cdot 10^{-31} kg" is the electon mass, "c = 3\cdot 10^8" speed of light in vacuum and "v" is the electron speed.

Finaly obtain:

"\lambda = \dfrac{h}{mv}\sqrt{1-\frac{v^2}{c^2}}" .

Substitute numerical values:

a) "\lambda = \dfrac{6.6\cdot 10^{-34}}{9.1\cdot 10^{-31}\cdot 1\cdot10^8}\sqrt{1-\frac{1\cdot10^{16}}{9\cdot 10^{16}}} = 6.84\cdot 10^{-12} m"

b) "\lambda = \dfrac{6.6\cdot 10^{-34}}{9.1\cdot 10^{-31}\cdot 2\cdot10^8}\sqrt{1-\frac{4\cdot10^{16}}{9\cdot 10^{16}}} = 2.7\cdot 10^{-12} m"